Help Seve &B 4 Calculate the four components of aggregate expenditure and GDP for the following economy using data from the table below Instructions: Enter your responses as whole numbers. If you are entering any in front of those numbers. numbers, be sure to include a 25 GDP 575 200 Construction of new homes and Print Imports $100 125 s1ee 5180 stocks Consurmption expenditures $ [ 60g Government Purchases $ Net Exports: $ 29 GDP: S Next >

GDP
Answer #1

Consumption Expenditure = Consumption Expenditures + Household purchase of durable goods = 600+150= $750

Investment Expenditures = Construction of new homes +Change in inventory stocks + Business fixed investments = 100+25+100 =$225

Government Purchases = $75

Net Exports = Export – Import = (75-50) =$25

GDP = Consumption +Investment + Government Purchases + Net Exports

GDP = 750+ 225+75+ 25 = $1075